OpenStudy (anonymous):

Find the identity element of (Q,*)? If (Q,*) is map to (Q,X) given by map(x)=3x-1 is group homorphism, where the binary operation in (Q,*) is a*b=3ab-a-b+2/3. Hint: when the binary operation is a*b=1/3(ab-a-b+2). The identity element is 2.

OpenStudy (anonymous):

i guess after teasing out all this weird verbiage, what you are being asked to do is find the inverse image of 2

OpenStudy (anonymous):

since the identity must go to the identity and vice versa

OpenStudy (anonymous):

no 2 is the answer for another question however i not sure of the working for either one

OpenStudy (anonymous):

oh ok. i am not sure what all that other business is about then, the homomphism etc

OpenStudy (anonymous):

i think you find this as follows you know \(a*e=a\) because \(e\) is the identity, and \[a*e=3ae-a-e+\frac{2}{3}\]by definition of \(*\) so set them equal and solve for \(e\)

OpenStudy (anonymous):

you get \[a=3ae-a-e+\frac{2}{3}\] or \[2a=3ae-e+\frac{2}{3}\] telling you \(e=\frac{2}{3}\)